OpenStudy (anonymous):

Let a and b be arbitrary vectors of constant length and angle, which are neither parallel nor perpendicular, and let c be the vector defined by the equation c=(cost)a + (sint)b. t is greater than or equal to 0. a) For what values of t is c parallel to a? b) When is c parallel to b? c) Can c ever be parallel to a + b? d) Can c ever be perpendicular to a + b?

OpenStudy (phi):

c will be parallel to a if it is c= M a where M is some number in other words, you don't want to add any of vector b. that means you want sin t = 0 (to make b "go away") what angle t has sin t = 0 ?

OpenStudy (anonymous):

Is it 0 or 180?

OpenStudy (phi):

yes, or in general n*180º where n is an integer or n* pi (in radians)

OpenStudy (phi):

a) For what values of t is c parallel to a? for t= n pi radians, n an integer

OpenStudy (anonymous):

So would the value possibly be pi?

OpenStudy (phi):

yes, pi works. sin(pi)=0 but sin(2pi= 0 sin(3pi)= 0 sin(-6 pi) = 0 and so on

OpenStudy (phi):

for b) When is c parallel to b? it's the same idea, except now you want cos t = 0

OpenStudy (anonymous):

So then you would want answers like 2pi/3 or 3pi/2?

OpenStudy (phi):

pi/2 or 3pi/2 (not 2 pi /3 ) you can test those in your calculator.

OpenStudy (anonymous):

Okay. Perfect. All of this makes so much more sense now. For questions c and d the answers would be no then because one of then has to be 0 in order for one of them to be parallel. Or is zero a possible answer or 180 as it was brought up earlier?

OpenStudy (phi):

c) Can c ever be parallel to a + b? to be parallel to a vector you have to be a scalar multiple of that vector for example, for c to be parallel to vector a we need c = M a where M is a number. if we can find an M so that c= M (a + b) then c will be parallel to (a+b) in other words, can we find c= M a + M b ? and that means can we find cos t = sin t (equal to the same number ?)

OpenStudy (phi):

start with the requirement sin t = cos t divide both sides by cos t \[ \frac{\sin(t)}{\cos(t)} = 1 \] or \[ \tan( t) = 1 \] what angles does tan t =1

OpenStudy (anonymous):

Okay. pi/4 and 5pi/4 ?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Perfect! This would be the answer then for c because of the one. Would d be any different?